OpenStudy (anonymous):

find two consecutive integers whos sum is 73

OpenStudy (nikato):

let the first integer be n the next consecutive integar would be n+1 and you know these 2 integers add up to 73 so n+(n+1)=73 n+n+1=73 combine like terms 2n+1=73 subtract 1 on both sides 2n=72 divide both sides by 2 so n=?

OpenStudy (anonymous):

by applying the above formula , we will get the consecutive terms to be 36 and 37 36+37=73